### v1

# DynaStride: Dynamic Stride Windowing with MMCoT for Instructional Multi-Scene Captioning

Abstract
Scene-level captioning in instructional videos can enhance learning by requiring an understanding of both visual cues and temporal structure. By aligning visual cues with textual guidance, this understanding supports procedural learning and multimodal reasoning, providing a richer context for skill acquisition. However, captions that fail to capture this structure may lack coherence and quality, which can create confusion and ultimately undermine the video’s educational intent. To address this gap, we introduce DynaStride, a pipeline to generate coherent, scene-level captions without requiring manual scene segmentation. Using the YouCookII dataset’s scene annotations, DynaStride performs adaptive frame sampling and multimodal windowing to capture key transitions within each scene. It then employs a multimodal chain-of-thought process to produce multiple action-object pairs, which are refined and fused using a dynamic stride window selection algorithm that adaptively balances temporal context and redundancy. The final scene-level caption thus integrates visual semantics and temporal reasoning in a single instructional caption. Empirical evaluations against strong baselines, including VLLaMA3 and GPT-4o, demonstrate consistent gains on both N-gram–based metrics (BLEU, METEOR) and semantic similarity measures (BERTScore, CLIPScore). Qualitative analyzes further show that DynaStride produces captions that are more temporally coherent and informative, suggesting a promising direction for improving AI-powered instructional content generation.
Introduction
Instructional videos are widely used to teach complex tasks by providing step-by-step guidance that can be modeled through deep learning (Narasimhan et al. 2023). These videos typically represent a sequence of steps, each with distinct scenes that highlight specific stages of the task. Recent work shows that instructional video corpora often contain rich multimodal signals which can be exploited to segment and caption procedural steps, but also introduce challenges of alignment, redundancy, and temporal coherence (Shi and Ji 2019). Generating accurate scene-level captions is crucial for accessibility, content summarization, and video search, enabling learners to quickly grasp sequences of actions and the underlying concepts (Yousif 2023). High-quality captions also improve accessibility for visually impaired or differently abled learners and allow educators to efficiently index and summarize content (Gernsbacher 2015). In addition, instructional videos have been shown to significantly improve learning outcomes, with a meta-analysis revealing a strong effect size of 0.688 when video clips are thoughtfully integrated into lessons (Lin and Yu 2023). As such, advances in scene-level caption generation have significant potential to enhance personalized learning experiences, enable adaptive recommendations, and support the delivery of scalable educational content.
Artificial Intelligence (AI) and Machine Learning (ML) have increasingly transformed educational technologies, allowing adaptive learning, personalized feedback, and data-driven insights into student performance (Elstad 2024; Morales-Navarro and Kafai 2024). In the context of instructional videos, AI-powered video analysis and automated captioning systems can improve accessibility, facilitate content indexing, and support personalized learning experiences (Shvetsova et al. 2024). Deep learning models, including vision-language architectures, allow for automated understanding of complex visual and temporal sequences, providing learners and educators with actionable insights from video content (Chu, Seo, and Han 2025; Yang et al. 2023). Together, these AI and ML advancements promise to make instructional video-based learning more accessible, engaging, and tailored to individual learner needs.
Recent research has highlighted the importance of scene-level captions in educational videos. For instance, Malakul and Park (2023) demonstrated that automatically generated subtitles can significantly enhance both learning comprehension and learner satisfaction compared to videos without subtitles. Advances in video captioning, such as leveraging scene graphs for fine-grained captioning, have further shown promise in producing captions that are more accurate and contextually aware, especially for longer videos (Chu, Seo, and Han 2025). Additionally, Alabsi (2020) found that using apps to add subtitles to educational videos can improve students’ listening comprehension, indicating that the way subtitles are presented can meaningfully affect learning outcomes (Alabsi 2020). These studies demonstrate that high-quality, scene-level captions can significantly improve how educational videos support learning and comprehension.
Despite significant progress in scene-level captioning, current methods face notable challenges in accurately representing instructional videos. Many approaches attempt to generate captions for every frame or segment, often producing dense descriptions that include redundant or irrelevant information (Chai et al. 2024; Yang et al. 2023). In contrast, shorter caption methods frequently miss important details about actions, objects, or their temporal relationships, resulting in incomplete or ambiguous descriptions (Tang et al. 2025). Furthermore, some systems rely heavily on localized inference tools, which can limit scalability, reproducibility, and the ability to maintain consistent quality in long or complex videos (Chai et al. 2024). Ultimately, these challenges point to the need for smart, context-sensitive captioning that captures key actions and objects, reduces repetition, and preserves the correct sequence of events.
In this work, we introduce DynaStride, a scalable framework for scene-level captioning on educational videos, evaluated on the YouCookII dataset (Zhou, Xu, and Corso 2018). By leveraging existing scene boundaries and human-labeled ground-truth captions, DynaStride avoids the need for custom segmentation. This framework adaptively skips redundant frames, applies multimodal chain-of-thought (MMCoT) prompting (Zhang et al. 2025) to generate multiple detailed subcaptions, and finally aggregates them into concise, scene-level descriptions. When evaluating DynaStride against state-of-the-art vision-language models, including GPT-4o (Hurst et al. 2024) and VideoLLaMA-3 (Grattafiori et al. 2024), our framework demonstrates improvements in both N-gram and semantic-based metrics.
Related Works
Video Captioning
Recent video captioning research has focused on improving temporal understanding and long-form video comprehension using encoder-decoder and transformer-based architectures that capture dense, action-oriented representations (Xie et al. 2024; Wu et al. 2025; Muksimova et al. 2025). Hierarchical approaches, such as AuroraCap and ShareGPT4Video (Chai et al. 2025; Chen et al. 2024), decompose the task into stages including feature extraction, event localization, and caption generation across temporal scales , while multi-level frameworks generate captions at the frame, clip, and video levels (Wei et al. 2025). Other methods enhance temporal reasoning with mechanisms like Timestamp Injection Mechanism (TIM) and Temporal-Aware Similarity Sampling (TASS) (Yuan 2025), or use causal-temporal narrative structures for coherent descriptions (Nadeem and Mustafa 2024). Reinforcement learning and intention-oriented controllable captioning improve fine-grained or targeted descriptions (Meng and Zhang 2025; Qiu and Gao 2025), and spatial-temporal and multimodal modules like STC enhance video-language modeling (Cheng and Leng 2024). Building on these strategies, DynaStride aggregates frame and scene-level captions, combining temporal reasoning with long-form video understanding.
Chain-of-Thought Reasoning
Wei et al. (2022) introduced Chain-of-Thought (CoT) prompting, a technique that enhances the reasoning capabilities of large language models by guiding them to generate intermediate reasoning steps before producing a final answer. This approach has been particularly effective for tasks requiring multi-step reasoning, such as arithmetic, commonsense, and symbolic reasoning, by encouraging models to mimic human-like thought processes. Building on this idea, Zhang and Li (2025) proposed the Complex Multi-Modal Chain-of-Thought (CMMCoT) framework, which extends CoT reasoning to complex multi-image comprehension tasks. CMMCoT employs interleaved multimodal reasoning chains, where critical visual region tokens extracted from intermediate reasoning steps serve as supervisory signals, improving cross-modal understanding and interpretability. Additionally, it incorporates test-time memory augmentation through the Retrieval-based Image Feature Reasoning Enhancement Module (RIFREM), which stores key and value pairs of image features from each decoder layer in a memory bank. This design enables efficient cross-attention computations and allows the model to dynamically recall relevant visual information during reasoning, effectively expanding its reasoning capacity without increasing the number of parameters. Collectively, these studies highlight the growing importance of structured reasoning and dynamic memory mechanisms for advancing multimodal reasoning across video frames.
Figure 1: Illustration of the multi-scene-to-text captioning pipeline depicting four stages: frame sampling, MMCoT, dynamic stride, and aggregation.
Methodology
To achieve our goal of generating high-quality instructional captions for multiple scenes, we propose the following multi-scene to text pipeline: (1) frame sampling and windowing, (2) MMCoT subcaption generation, (3) a dynamic stride window selection algorithm, and (4) subcaption aggregation.
The goal of this pipeline is to reduce content redundancy in scenes by eliminating semantically similar windows and retaining only the most informative ones. This allows MMCoT to more effectively track actions over time and reason about the objects involved. Pretrained models were used without fine-tuning to ensure the reproducibility of our pipeline.
Frame Sampling and Windowing
To efficiently process long video scenes, we first define the $i$-th scene of a video $V$ as an ordered sequence of frames:
$V_{i}=[f_{i1},f_{i2},\dots,f_{iN_{i}}],\quad f_{ik}\in\mathbb{R}^{H\times W\times C},\quad 1\leq k\leq N_{i},$
where $N_{i}$ is the total number of frames in the scene, and $H,W,C$ denote the height, width, and number of color channels of each frame, respectively.
Directly processing all frames in a long scene can be computationally expensive and often redundant, as consecutive frames typically contain highly similar visual information. To address this, we subsample the scene by selecting every $M$-th frame, producing a more compact subsequence:
$\tilde{V}_{i}=[\tilde{f}_{i1},\tilde{f}_{i2},\dots,\tilde{f}_{i\lfloor N_{i}/M\rfloor}].$
Next, to capture localized temporal dynamics while keeping computations manageable, we define a sliding window $W_{it}$ starting at index $t$ as a consecutive block of $K$ frames from the subsampled sequence:
$W_{it}=[\tilde{f}_{it},\tilde{f}_{i(t+1)},\dots,\tilde{f}_{i(t+K)}].$
This windowing strategy allows our model to focus on short-term temporal patterns within each scene, facilitating more precise feature extraction and downstream caption generation, while avoiding unnecessary processing of redundant frames.
MMCoT Subcaption Generation
For each window $W_{it}$, we first concatenate its frames horizontally to form a wide image:
$X_{it}=\text{Concat}(W_{it})\in\mathbb{R}^{H\times(|W_{it}|\cdot W)\times C},$
which allows a vision-language model to perceive temporal context across multiple frames in a single input, rather than processing each frame independently.
We then employ Qwen3 ($\mathcal{Q}$) (Team 2024), an image-to-text model, to generate a subcaption describing both the observed action and associated objects:
$c_{it}=\mathcal{Q}(X_{it})=\text{"$a_{it}\,|\,o_{it}$"},$
where $a_{it}$ represents the action description, $o_{it}$ is the set of objects present, and the delimiter “||” separates these components to facilitate downstream aggregation. This multimodal chain-of-thought approach encourages the model to reason over both temporal and semantic content simultaneously.
To reduce redundancy and improve efficiency, we adaptively select windows using a stride parameter $s$. Let $p$ denote the index of the most recently retained window with subcaption $c_{ip}$. For a candidate window $W_{it}$, we first compute its embedding using MiniLM ($\mathcal{M}$) (Wang et al. 2020):
$e_{it}=\mathcal{M}(c_{it}),\quad e_{ip}=\mathcal{M}(c_{ip}),$
and measure similarity via cosine similarity: $\text{sim}(e_{it},e_{ip})$.
If $\text{sim}(e_{it},e_{ip})\geq\tau$, the candidate window is skipped, and the stride is increased according to
$s\leftarrow\min(\alpha\cdot s,s_{\max}),$
Dynamic Stride Window Selection.
Algorithm 1 Dynamic Stride Window Selection
Input: Video frames $V_{i}$, window size $K$, subsample rate $M$
Parameter: Base stride $s_{\text{base}}$, max stride $s_{\max}$, stride scaling factor $\alpha$, similarity threshold $\tau$
Output: Video captions $\mathrm{captions}$
[…]
6:while $t<|\tilde{V}_{i}|$ do
7:  $W_{it}\leftarrow\tilde{V}_{i}[t:t+K]$
8:  $X_{it}\leftarrow\operatorname{Concat}(W_{it})$
9:  $c_{it}\leftarrow\mathcal{Q}(X_{it})$
10:  $e_{it}\leftarrow\mathcal{M}(c_{it})$
11:  if $\operatorname{sim}(e_{it},e_{ip})\geq\tau$ then
12:   $s\leftarrow\min(\alpha\cdot s,s_{\max})$
13:  else
14:   $\mathrm{captions}$.append($c_{it}$)
15:   $e_{ip}\leftarrow e_{it}$
16:   $s\leftarrow s_{\text{base}}$
17:  end if
18:  $t\leftarrow t+\max(1,s)$
19:end while
20:return $\mathrm{captions}$
where $\alpha\geq 1$ is a scaling factor and $s_{\max}$ is the maximum allowable stride. Otherwise, the window $W_{it}$ is retained, its subcaption $c_{it}$ is selected, and the stride resets to the base value $s\leftarrow s_{\text{base}}$.
Algorithm 1 provides a detailed description of this procedure, which dynamically balances coverage of temporal events with computational efficiency, ensuring each subcaption contributes meaningful information to the final caption.
Subcaption Aggregation
The retained subcaptions $\{c_{i1}^{\prime},\dots,c_{iL_{i}}^{\prime}\}$ are concatenated and fed into Qwen2.5 ($\mathcal{A}$) to produce a final instructional caption for the scene:
$C_{i}=\mathcal{A}(c_{i1}^{\prime}\oplus\cdots\oplus c_{iL_{i}}^{\prime}),$
[…]
Implementation details, including parameter choices for $M$, $K$, $\alpha$, $s_{\text{base}}$, and $\tau$, are provided in the Appendix.
Experiments
In this section, we present a series of empirical experiments designed to evaluate the performance of DynaStride and investigate its potential limitations. Our goal is to gain a deeper understanding of how DynaStride handles scene-level captioning and the factors that influence its effectiveness. We structure our investigation around the following research questions:
Method N-gram BERT SBERT
B@4(↑) METEOR(↑) CIDEr(↑) Prec(↑) Recall(↑) F1(↑) Sim(↑)
GPT-4o 4.73 (0.63) 28.47 (1.37) 0.48 (0.06) 0.19 (0.01) 0.29 (0.01) 0.23 (0.01) 0.60 (0.01)
VLLaMA-3 4.10 (0.32) 22.71 (0.63) 0.49 (0.02) 0.19 (0.01) 0.22 (0.00) 0.21 (0.01) 0.58 (0.00)
DynaStride 4.18 (0.07) 24.31 (0.10) 0.56 (0.00) 0.25 (0.00) 0.26 (0.00) 0.27 (0.00) 0.61 (0.00)
Table 1: Scene captioning results on YouCook2 validation set, comparing GPT-4o, VLLaMA-3, and DynaStride. Best results per metric are in bold.
  • RQ 1. To what extent does our method improve the coherence and meaningfulness of the inferred scene captions?
  • RQ 2. How does the frame sampling and aggregator impact overall caption quality?
Dataset
We evaluated our method on the widely used YouCookII dataset (Zhou, Xu, and Corso 2018), which contains 2,000 long-form cooking videos, each annotated with scene boundaries and their corresponding instructional descriptions. As the official test set captions are not publicly available, we uniformly sampled 210 videos from the validation set for all experiments.
To prepare the dataset for downstream scene captioning tasks, we implemented a preprocessing pipeline that converts raw videos and annotations into scene-segmented, frame-level data suitable for model training and evaluation. Starting from the official YouCookII dataset archive, which includes raw videos, JSON annotations, and split lists, we organized the data by split and recipe type to facilitate batch processing. We did not utilize the training split as our pipeline does not require training.
For reliable video retrieval, we replaced the original youtube-dl with yt-dlp and handled restricted videos using manual cookie export and authentication. All videos were saved in .mp4 format and verified for integrity.
Annotations were parsed to extract scene boundaries (startFrame, endFrame) and stored as .pkl files. Videos were segmented according to these boundaries, and frames within each scene were sampled and resized to $\times384$ pixels to standardize model inputs. Scene frames were stored in uniquely named directories (<video_id>_<segment_num>) under their respective split folders, allowing efficient access during training and evaluation.
Experimental and Ablation Setup
We evaluated DynaStride against two SoTA VLMs, GPT-4o (Hurst et al. 2024) and VideoLLaMA-3 (Grattafiori et al. 2024), using the same prompt for both baselines and the aggregator model to ensure fair comparison. Experiments and ablations were conducted on three seeds, reporting mean and standard deviation for all metrics. For computing, we used 30× RTX A6000 GPUs and 800 GB of storage. Performance was measured across multiple captioning quality metrics covering N-gram overlap, semantic alignment, TF-IDF consensus, and embedding similarity: BLEU-4 (Papineni et al. 2002), METEOR (Banerjee and Lavie 2005), CIDEr (Vedantam, Zitnick, and Parikh 2015), BERTScore (Zhang et al. 2020) (Precision, Recall, F1), SBERT Similarity (Reimers and Gurevych 2019), and Temporal Analysis (Sun et al. 2021; Ilharco et al. 2019) (see Appendix for details). We also investigated the impact of frame sampling rates and aggregator choice on caption quality, comparing sampling rates (5, 20, 40) and aggregators including Phi (Abdin et al. 2024) and Mistral (Jiang et al. 2023). Direct caption examples and graphs of frame sampling rates versus performance metrics are in the Appendix (Figure 2 and 3).
DynaStride achieves the highest CIDEr score and semantic metrics compared to the baselines
As shown in Table 1, DynaStride consistently outperforms VLLaMA-3 across all metrics, with notable gains in CIDEr, precision, F1, and SBERT similarity. Compared to GPT-4o, DynaStride achieves higher scores on CIDEr, BERT precision, BERT F1, and SBERT similarity, indicating that our captions are not only more relevant but also semantically closer to the ground-truth references. These results highlight that our method produces captions that are both more accurate and contextually meaningful. Further results on temporal performance, which measures how accurately the model captures the timing and order of events within a video, can be found in the Appendix section (Table 3-5).
Configurations N-gram BERT SBERT
B@4(↑) METEOR(↑) CIDEr(↑) Prec(↑) Recall(↑) F1(↑) Sim(↑)
Aggregator Comparison
   Phi 2.78 (0.24) 22.8 (0.38) 0.37 (0.02) 0.17 (0.01) 0.25 (0.00) 0.21 (0.01) 0.59 (0.00)
   Mistral 3.36 (0.08) 19.49 (0.12) 0.51 (0.01) 0.27 (0.00) 0.23 (0.00) 0.25 (0.00) 0.60 (0.00)
   Qwen3 4.18 (0.07) 24.31 (0.10) 0.56 (0.00) 0.25 (0.00) 0.26 (0.00) 0.27 (0.00) 0.61 (0.00)
Frame Sampling Rates
GPT-4o
   5 Frames 4.31 (0.15) 27.58 (0.47) 0.45 (0.02) 0.18 (0.00) 0.28 (0.00) 0.23 (0.00) 0.59 (0.00)
   20 Frames 4.69 (0.19) 27.91 (0.28) 0.49 (0.02) 0.19 (0.01) 0.29 (0.00) 0.24 (0.00) 0.60 (0.00)
   40 Frames 4.52 (0.11) 28.07 (0.03) 0.48 (0.00) 0.19 (0.00) 0.29 (0.00) 0.24 (0.00) 0.60 (0.00)
VLLaMA-3
   5 Frames 3.60 (0.45) 22.48 (0.17) 0.45 (0.05) 0.18 (0.00) 0.22 (0.00) 0.19 (0.00) 0.57 (0.00)
   20 Frames 4.32 (0.29) 22.28 (0.27) 0.52 (0.02) 0.22 (0.00) 0.22 (0.00) 0.22 (0.00) 0.58 (0.00)
   40 Frames 4.79 (0.05) 21.90 (0.08) 0.56 (0.01) 0.27 (0.00) 0.21 (0.00) 0.24 (0.00) 0.58 (0.00)
DynaStride
   5 Frames 3.89 (0.12) 23.38 (0.14) 0.52 (0.01) 0.24 (0.00) 0.25 (0.00) 0.24 (0.00) 0.60 (0.00)
   20 Frames 4.48 (0.03) 24.82 (0.10) 0.58 (0.00) 0.24 (0.00) 0.26 (0.00) 0.25 (0.00) 0.61 (0.00)
   40 Frames 4.91 (0.03) 26.36 (0.18) 0.61 (0.00) 0.25 (0.00) 0.28 (0.00) 0.27 (0.00) 0.63 (0.00)
Table 2: Ablation study comparing aggregator models (Phi, Mistral, Qwen3) and frame sampling rates (5, 20, 40 frames) across GPT-4o, VLLaMA-3, and DynaStride. Metrics include N-gram (BLEU-4, METEOR, CIDEr), BERT-based (Precision, Recall, F1), and SBERT similarity. Bold values indicate the best performance per metric.
Sparse sampling boosts caption quality and aggregator choice impacts consistency
Based on the metrics in Table 2, higher frame sampling generally improves caption quality across all models, though the effect varies by method. For GPT-4o, increasing from 5 to 20 frames yields consistent gains in B@4, METEOR, and CIDEr, plateauing between 20 and 40 frames. VLLaMA-3 shows steady improvement up to 40 frames, achieving the highest CIDEr and F1 scores, although SBERT similarity remains lower than that of GPT-4o. DynaStride benefits the most from sparser sampling, with 20 and 40 frames producing consistent gains across all metrics, reaching the highest overall scores at 40 frames; particularly CIDEr (0.61), F1 (0.27), and SBERT similarity (0.63). We also see aggregator model choice has a strong impact on captioning consistency, particularly on N-gram metrics. Qwen3 dominates overall performance in all but BERT Precision, and has the lowest standard deviation across all metrics. Furthermore, we found that Phi’s standard deviation metrics were more than three times higher than both Qwen3 and Mistral, indicating that it had the least consistent caption generations. These results indicate that both dense frame sampling and careful aggregator selection are crucial. Dense sampling helps produce coherent and semantically meaningful captions, while the choice of aggregator strongly influences consistency.
Limitations
Despite its promising performance, DynaStride has several limitations that are important to consider. Its reliance on pretrained models such as Qwen2.5-VL and Qwen3 raises concerns about generalization beyond the YouCookII domain, as these models may not fully capture domain-specific nuances or procedural variations in other instructional contexts. Similarly, the YouCookII dataset, while widely used, is relatively small and may not reflect the full diversity of instructional tasks, which could limit the applicability of the generated captions to other domains or more complex procedural workflows.
Additionally, the current architecture does not explicitly model cross-modal interactions over long temporal spans, which may reduce its ability to capture complex dependencies between actions, objects, and instructional narration. While dynamic frame sampling improves computational efficiency, it may overlook subtle or rapid actions, potentially producing incomplete, ambiguous, or temporally inconsistent subcaptions. Although the dynamic stride algorithm reduces redundant frames, it may not completely mitigate this issue during subcaption aggregation, potentially impacting the coherence, clarity, and instructional value of the final scene-level captions.
Moreover, standard automatic evaluation metrics, such as N-gram overlap or semantic similarity scores, fail to fully capture human judgments of instructional relevance, temporal accuracy, or educational usefulness, which are crucial for learning outcomes. Finally, DynaStride does not currently incorporate mechanisms for domain adaptation or active feedback, limiting its ability to continuously improve in new instructional settings or personalize captions for diverse learner populations.
Conclusion
This work introduces DynaStride, a pipeline that leverages YouCookII annotations for scene-level captioning in instructional videos. By applying frame sampling to focus on the most informative portions of each scene, generating MMCoT action-object subcation pairs with Qwen2.5-VL, filtering redundant pairs with the dynamic windowing algorithm, and aggregating them into coherent scene-level captions with Qwen3, DynaStride efficiently captures key instructional content. Thus, DynaStride is able to support educational tasks such as lesson indexing, content retrieval, comprehension, and accessibility, aligning with broader efforts to enhance learning through AI-powered instructional content generation.
DynaStride offers several practical benefits. Dynamic frame sampling reduces computational overhead while retaining essential visual information, allowing faster processing without major sacrifices in caption quality. Aggregation of subcaptions ensures that scene-level captions remain coherent, informative, and temporally consistent, effectively conveying sequences of actions critical for understanding procedural content. Quantitative evaluation shows that DynaStride outperforms baseline models across N-gram and semantic metrics, including CIDEr, precision, F1, and SBERT similarity, indicating strong performance in both lexical accuracy and semantic understanding.
Despite these strengths, several limitations remain, including reliance on pretrained models, limited domain diversity in YouCookII, and challenges in modeling long-range cross-modal dependencies. Addressing these limitations represents a promising direction for future work. Extending the pipeline to handle raw, unsegmented videos will require robust scene boundary detection, potentially through temporal action detection models or weakly supervised segmentation techniques. Adaptive frame sampling informed by visual saliency, action dynamics, or instructional importance could further optimize coverage while maintaining efficiency. Improved caption aggregation that models procedural coherence and temporal alignment could enhance instructional clarity. Finally, incorporating human evaluations and expanding to diverse instructional domains would ensure that captions are both practically useful and educationally impactful.